INAF queue observing, 2016 Jan 25-26

Observer: F. Cusano, A. Rossi
Telescope Operator: David Gonzalo-Huerta
Support Astronomer: Barry Rothberg
Instrument: LUCI/MODS

Summary

At the sunset the sky is clear.

Night duration (12\xB0 el twilight) 19:20-6:10 ( 02:20-13:10 UT) = 10 h 50m

Observing time: 8 h 30m (85%)

Weather time loss: 1 h 0m (10%)

technical time loss 0.5 h 0m (5%)

Data Summary

Project_RUN ObjectsSorted descending Filter ExpTime Notes
Antoniucci V2775 Ori dual grating 0.3h MODS
Antoniucci V2775 Ori 200 H+K grating,zJspec, HKspec 0.3h LUCI only part of 1 script
Antoniucci V1647 Ori dual grating 0.3h MODS
Antoniucci V1118 Ori dual grating 0.3h MODS
Flat LUCI skyflat J, Ks, FeII, H2    
DeBarros MACS1423 200 H+K 0 min wrong allignement, nothing useful
Guaita M1149 H2 0.5h passing clouds!
Guaita M0717 H2 0.5h  
Sarracco M004XLSms2 red-grating 1h check if data are within constraints
Gilli id3605 200 H+K grating,zJspec 1.5h  
Darks     5x30,5x22,5x180,5x120, 3x300 for Antoniucci, Gilli, DeBarros, Polletta

Details

00:00 We start dark with LUCI
  • Gilli and Debarros dark luci2.20160126.0001-0003
  • Antoniucci dark luci2.20160126.0004-0023 (missing the 150s dark we do it tomorrow)

01:00 staring LUCI sky flat with N3.75 cam
  • J luci2.20160126.0025-0032 (24 not good guider in the field)
  • Ks luci2.20160126.0033-0039
  • FeII luci2.20160126.0040-0041 (counts too low)
  • J luci2.20160126.0042-0054 (42 low counts)

01:58 we start with MODS, slew to Saracco field M004XLSms2
  • acqMODS mods.4.XLSms2.acq mods1r.20160126.0002-0004
  • execMODS mods.4.XLSms2_5.obs mods1r.20160126.0005-0008
  • seeing 1.2" at 2:14
  • seeing 1.4" at 2:40
  • seeing 1.6" at 3:00

03:22 slew to standard GD71
  • acquisition mods1r.20160126.0009-0010
  • spectra mods1r.20160126.0011-0013 red grating
  • seeing 1.5"

3:45 the seeing is 1.6" we move to Antoniucci mods program, we start doing the GD71 in dual grating (using the previous acquisition)
  • spectra mods1r.20160126.0014-0015, mods1b.20160126.0002-0003 dual grating

3:58 slewing to V1118 Ori
  • acquisition mods1r.20160126.0016-0018
  • seeing 1.8" at 4:06
  • spectra mods1r.20160126.0019-0022, mods1b.20160126.0004-0007
  • seeing 1.2" at 4:19
  • seeing 2" at 4:26

04:36 slew to V1647 Ori
  • the allskycamera is frozen, there is 25% humidity but in the dome 90%, the TO is checking, the temperature changed 4 degrees in 10 min
  • acquisition (wrong guide star, we changed) mods1r.20160126.0023-0025
  • spectra mods1r.20160126.0026-0029, mods1b.20160126.0008-0011
  • seeing 1.2" at 4:55, seein 1.6" after two minutes, seeing very variable and temperature gradient
  • seeing 1.6" at 5:22, temperature dropping down and humidity climbing we stay on the Antoniucci program

5:23 slew to V2775 Ori
  • acquisition mods1r.20160126.0030-0032
  • spectra mods1r.20160126.0033-0036, mods1b.20160126.0012-0015

06:08 we change to LUCI2

06:10 we are going to check the condition of the seeing on the standard P-161D for Guaita
  • we slew to P-161D, luci2.20160126.0055-0084 J, Br_Gam, H2
  • seeing 1" on image 6:25, seeing variable during the exposure
  • seeing 0.9" on image 6.48"

06:50 we slew to Guaita H2 filter
  • problem is not possible to reach collimation, the seeing is too variable
  • we change target

07:11 moving to Antoniucci program, object V2775 Ori with LUCI
  • problem in finding the guide star, the TO do the collimation again
  • acq luci2.20160126.0085-0088
  • seeing 0.9" at 7:23 grrrrrrr
  • ahhh seeing 1.5" at 7:29
  • spectra luci2.20160126.0089-0093 we toke three spectra in zJ and two in HJ because the object is too low and we are reaching the limit

07:56 slew to the telluric of V2774 Ori
  • scheduler stucked several time during the acquistion
  • acq luci2.20160126.0094-0099
  • spec luci2.20160126.0100-0103
  • seeing stable to 1" since 15 minutes we move to Guaita program

08:20 we slew to Guaita H2
  • luci2.20160126.0104-133
  • checking pointing 8:23
  • seeing 0.8" on the guider, 0.7" on the images
  • seeing 0.5" on the images at 08:32
  • seeing 0.7" on the images at 09:08

09:13 we slew to Gilli telluric
  • acq luci2.20160126.0134-0136
  • spec luci2.20160126.0137-0139
  • seeing 0.8" on the guider

09:26 we slew to Gilli id3605
  • acq luci2.20160126.0140-0143
  • spec luci2.20160126.0144-0161
  • seeing 0.8" on the guider at 9:37
  • seeing 0.6" on the guider at 10:12
  • thin cirrus coming in at 10:30
  • seeing 0.5" on the guider at 10:45

11:20 we slew to deBarros telluric
  • telluric acquisition check pointing;
  • problem with the acquisition the images are not showing in the longslit alignment panel,
  • acq luci2.20160126.0162-0165
  • spec luci2.20160126.0166-174

11:53 we slew to deBarros science
  • check pointing
  • seeing 0.6" at 12:02 on the guider
  • acq luci2.20160126.0175-184
  • problem with the acquisition due to a scheduler problem
  • started again but passing clouds did not allow us to complete the acquisition

12:40 we slew to Guaita_CD
  • luci2.20160126.0185-0214
  • check pointing
  • seeing 0.6" at 12:50 on the guider
  • background affected by passing clouds, use with caution
  • seeing 0.5" at 12:58 on image

13:40 sky flat with luci
  • J luci2.20160126.0215-0223
  • Ks luci2.20160126.0224-02
  • FeII 238-248
  • H2 249-263

14:25 calibrations for MODS, program Sarracco
  • mods1r.20160126.0037-00??

-- %USERSIG{Felice Cusano, Andrea Rossi - 2016-01-25}%

Comments

 
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback